OpenStudy (purple_pink):

A recipe that makes 7 servings calls for 1 1/6 cups of juice. How many cups of juice are there per serving? A. 1/6 B. 1/7 C. 7/6 D. 6/7

Vocaloid (vocaloid):

divide (1 1/6)/7

OpenStudy (purple_pink):

0.66666666... @Vocaloid

OpenStudy (purple_pink):

ohhhhh so it's A

Vocaloid (vocaloid):

check your math again

OpenStudy (purple_pink):

umm i used a calculator

Vocaloid (vocaloid):

(1+1/6)/7 = ?

OpenStudy (purple_pink):

2/6/7

Vocaloid (vocaloid):

nope... 1 + 1/6 = ?

OpenStudy (purple_pink):

3

Vocaloid (vocaloid):

uh, not quite, check your math again 1 + 1/6 = ?

OpenStudy (purple_pink):

2/6 @Vocaloid

Vocaloid (vocaloid):

remember how to add fractions 1 = 6/6 6/6 + 1/6 = ?

OpenStudy (purple_pink):

1 1/6

Vocaloid (vocaloid):

as an improper fraction

OpenStudy (purple_pink):

7/6

Vocaloid (vocaloid):

good, now (7/6)/7 = ?

OpenStudy (purple_pink):

0.166666...

OpenStudy (purple_pink):

@Vocaloid

Vocaloid (vocaloid):

as a fraction, please

OpenStudy (purple_pink):

1 1/6

OpenStudy (purple_pink):

@Vocaloid

Vocaloid (vocaloid):

nope, not quite